OpenStudy (anonymous):

Factor. x^2 + 7x + 10 A. (x + 2)(x + 5) B. (x + 1(x + 10) C. (x + 1)(x + 9) D. Prime

OpenStudy (anonymous):

Since you want numbers that add to 7, and multiply to 10, you can start by making a list of things that multiply to 10. 10*1 5*2 2*5 1*10 and then check adding 10 + 1 = 11 5+2=7 ----> this works !!!

OpenStudy (anonymous):

So we get (x+___)(x+___)

OpenStudy (anonymous):

does this make sense?

OpenStudy (anonymous):

To see the multi, and add thing.... expand \[(x+a)(x+b) = x^2+(a+b)x+ab\]

OpenStudy (anonymous):

@Redcan so which answer choice is that?